# Invariant Subspaces of Perturbed Backward Shift

### v1

@@ Abstract.
We represent closed subspaces of the Hardy space that are invariant under finite-rank perturbations of the backward shift. We apply this to classify almost invariant subspaces of the backward shift and represent closed subspaces that are invariant under a more refined version of nearly invariant subspaces of the backward shift. Kernels of certain perturbed Toeplitz operators are examples of the newly introduced nearly invariant subspaces.
@@
Hardy space, invariant subspace, shift operator, nearly invariant subspace, model space, Schmidt subspace
@@ 1. Introduction
In this paper, we bring together three problems concerning invariant subspaces. These problems have one commonality: the backward shift operator $T_{z}^{*}$ acting on $H^{2}$ appears in all of them. Recall the standard notation that $T_{z}$ is the multiplication operator by the coordinate function $z$; alternatively, the Toeplitz operator with the symbol $z$; and $H^{2}$ is the Hardy space of all square summable analytic functions on the open unit disc $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$ (see (2.3) for more details). Therefore
@@
for all $f\in H^{2}$. We effectively connect invariant subspaces of three types of operators: finite rank perturbations of $T_{z}^{*}$, nearly invariant subspaces of $T_{z}^{*}$ (in a larger sense), and almost invariant subspaces of $T_{z}^{*}$. Strategically, we first represent closed invariant subspaces of finite rank perturbations of $T_{z}^{*}$, and then we use that technique to represent and classify the invariant subspaces of the two subsequent classes of operators.
Therefore, the broad topic of this paper is invariant subspaces of $T_{z}^{*}$-related operators. Linear analysis has been contending with invariant subspaces from the very beginning. The famous “invariant subspace problem” calls for finding a nontrivial closed subspace that is invariant under a bounded linear operator on a Hilbert space. This problem is still open. Even worse, the existence of an invariant subspace for rank-one perturbations of diagonal operators is unknown [5]. In the same vein, the lattice of the invariant subspaces of simple composition operators is regarded as a difficult problem and is inextricably linked to the original invariant subspace problem [11]. Hence, identifying the lattices of invariant subspaces of natural operators is commonly a complex problem, and its solution often provides a valuable understanding of various topics (cf. Beurling theorem [1] and it’s applications [12, 13]). In this paper, we will also address all of these major issues related to $T_{z}^{*}$.
Before moving on, it is worth remarking that the results and approach of this paper are an enhancement of the circles of ideas that were presented by Hitt [9], which were further fine-tuned by Sarason. In fact, we identify Sarason’s observation on a specific rank-one projection [15, page 482] (also see (1.2) for more details) and use that perspective to provide full-length parametrizations of invariant subspaces of all finite-rank perturbations of $T^{*}_{z}$. We have subsequently laid out the road map to link the other two ideas.
@@
for all $f\in H^{2}$. Similarly, rank-$m$ operators are those whose ranges are $m$-dimensional. They can be represented as:
@@
where $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ are orthonormal sets of vectors in $H^{2}$. For the remainder of this section, we fix these orthonormal sets. We are interested in the invariant subspaces of a general finite-rank perturbation $X$ of $T_{z}^{*}$, where
@@
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $T_{z}^{*}+\sum_{i=1}^{m}v_{i}\otimes u_{i}$, then there exist $p^{\prime}\leq p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ such that
@@
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F\|^{2}+\|f_{0}\|^{2}=\|F_{0}F+f_{0}\|^{2}\}.$
In the given context, $H^{2}_{\mathbb{C}^{p}}$ represents the $\mathbb{C}^{p}$-valued Hardy space over $\mathbb{D}$, which may also be seen as $p$-copies of $H^{2}$. Given Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$, we denote $\mathcal{B}(\mathcal{H},\mathcal{K})$ as the space of all bounded linear operators (and write $\mathcal{B}(\mathcal{H})$ if $\mathcal{K}=\mathcal{H}$). The symbol $H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ denotes the set of $\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})$-valued bounded analytic functions on $\mathbb{D}$, where a function $\Phi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ is inner if
$\Phi(z)^{*}\Phi(z)=I_{\mathbb{C}^{p^{\prime}}}\qquad(z\in\mathbb{T}\;a.e.)$
Moreover, $[F_{0},1]_{1\times(p+1)}=[f_{1},\ldots,f_{p},1]$ is a row vector, and $\mathcal{K}_{\Psi}$’s elements are columns.
In fact, we prove Theorem 1.1 in a much higher generality (see Theorem 2.1): we simply replace the backward shift $T_{z}^{*}$ by the adjoint of an analytic Toeplitz operator $T_{\varphi}$, where the symbol $\varphi\in H^{\infty}$ is an inner function vanishing at the origin.
Before proceeding, we need the definition of nearly $T_{z}^{*}$-invariant subspaces (see Hayashi [8], Hitt [9], and Sarason [15]). A closed subspace $\mathcal{M}$ of $H^{2}$ is nearly $T_{z}^{*}$-invariant if
@@
for all $f\in\mathcal{M}\cap zH^{2}$. Now, we will provide a few comments regarding Theorem 1.1. First, this result deals with linear perturbations, a theory in and of itself that is well-known to be challenging (see the classic [10]). Second, this result parameterizes all of the invariant subspaces of finite rank perturbations of $T_{z}^{*}$, which was known to be a challenging goal. As far as we know, this is the first result in this direction. Third, we provide this result in the tradition of Hitt and Sarason’s classical argumentation. Indeed, Sarason showed in [15, page 481] that if $\mathcal{M}$ is a nearly $T_{z}^{*}$-invariant subspace then there is a unique function
@@
such that $\mathcal{M}$ is invariant under rank-one perturbation $T_{z}^{*}-T_{z}^{*}g\otimes g$, that is
@@
Sarason’s observation motivated us to attempt full-length representations of invariant subspaces of finite-rank perturbations, leading to the formation of Theorem 1.1. At the very start of this section, this point was briefly mentioned. Finally, we recall the representation of nearly $T_{z}^{*}$-invariant subspaces, which is attributed to Hitt [9] (also see Sarason [15]). Every nearly $T_{z}^{*}$-invariant subspace $\mathcal{M}\subseteq H^{2}$ admits the representation
@@
where $g$ is a unique function as in (1.1). Undoubtedly, the result in Theorem 1.1 is in line with the same fundamentals. As previously said, in the present case, we have effectively applied the core ideas that were first put forward by Hitt and Sarason a few decades ago.
Now we turn to almost invariant subspaces for $T_{z}^{*}$. In the setting of the backward shift operator, this was introduced by Chalendar, Gallardo-Gutiérrez, and Partington [4]. Let $T\in\mathcal{B}(\mathcal{H})$, and let $\mathcal{M}\subseteq\mathcal{H}$ be a closed subspace. We say that $\mathcal{M}$ is almost invariant for $T$ if there exists a finite-dimensional subspace $\mathcal{F}\subseteq\mathcal{H}$ such that
@@
The defect of the space $\mathcal{M}$ is the lowest possible dimension of such a space $\mathcal{F}$. Almost invariant subspaces for $T_{z}^{*}$ were classified in [4, Corollary 3.4] in terms of nearly invariant subspaces with finite defect indices. We also study the same thing, but we take a completely different approach: we enter into this problem through finite-rank perturbations of $T_{z}^{*}$. Indeed, a special version of Theorem 5.1 implies the following link: almost invariant subspaces and invariant subspaces of finite rank perturbations of the backward shift operator.
@@
Let $\mathcal{M}$ be a closed subspace of $H^{2}$, and let $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ be orthonormal sets in $H^{2}$. If $\mathcal{M}$ is invariant under
@@
then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect dimension at most $m$. Conversely, if $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect dimension $m$, then $\mathcal{M}$ is invariant under
@@
where $\{f_{i}:1\leq i\leq m\}$ is an orthonormal basis of the $m$ dimensional defect space.
This yields a complete characterization of almost invariant subspaces of $T_{z}^{*}$ (see Corollary 5.2):
@@
Let $\mathcal{M}\subseteq H^{2}$ be a closed subspace. If $\mathcal{M}$ is almost invariant under $T_{z}^{*}$, then there exist natural numbers $p^{\prime}<p$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
@@
where $F_{0}=[f_{1},\ldots,f_{p}]$, and $\{f_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}(u_{i}):i=1,\ldots,m\}$, and $\{u_{i}\}_{i=1}^{m}$ is an orthonormal basis for the defect space. Moreover
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|{F}\|^{2}+\|{f_{0}}\|^{2}=\|F_{0}F+f_{0}\|^{2}\},$ (1.4)
@@
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$, defines a unitary operator $\mathcal{J}:\mathcal{K}_{\Psi}\to\mathcal{M}$. Conversely, if $\mathcal{M}$ has the representation (1.3) for some orthonormal basis $\{f_{i}\}_{i=1}^{p}$ for $\mathcal{W}$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p})}$ with $\mathcal{K}_{\Psi}$ as in (1.4), and a unitary $\mathcal{J}$ as in (1.5), then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$.
It should be noted that Chalendar, Gallardo-Gutiérrez, and Partington were the first to classify almost invariant subspaces under $T_{z}^{*}$ [4, Corollary 3.4]. However, the above result differs significantly from [4], which compares nearly $T_{z}^{*}$-invariant subspaces of finite defect indices with almost invariant subspaces. Here, we connect them with finite rank perturbations of $T_{z}^{*}$. This, in particular, yields more concrete representations of the backward shift-invariant subspace $\mathcal{K}_{\Psi}$. See the end of Section 5 for more comments.
Now we turn to a more refined version of nearly $T_{z}^{*}$-invariant subspaces. The motivation for this arises partly from the matching theory of finite-rank perturbations of $T_{\varphi}^{*}$ as well as the potential of the theory of kernels of finite-rank perturbations of Toeplitz operators, which we will elaborate on shortly. We start by setting up the notations. Let $\{z_{k}\}_{k=1}^{n}$ be a sequence of points (possibly repeated) in $\mathbb{D}$, and assume that
@@
Evidently, nearly $T^{*}_{z,B_{1}}$-invariant subspaces are precisely nearly $T_{z}^{*}$-invariant subspaces. In Theorem 6.3, we present a complete description of nearly $T^{*}_{z,B_{n}}$-invariant subspaces of $H^{2}$:
@@
Let $\mathcal{M}$ be a nontrivial closed subspace of $H^{2}$. Then $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant if and only if there exist natural numbers $r^{\prime}\leq r\leq n$ and an inner function $\Theta\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{r^{\prime}},\mathbb{C}^{r})}$ such that $\Theta(0)=0$ and
@@
where $\{g_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B{H}^{2})$, and
@@
Kernels of Toeplitz operators are crucial in the context of examples of nearly $T_{z}^{*}$-invariant subspaces. In a comparable manner, we provide examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces. Curiously, certain dedicated finite-rank perturbations of Toeplitz operators result in nontrivial examples of $T^{*}_{z,B_{n}}$-invariant subspaces. More specifically, given the symbol $\varphi\in L^{\infty}$ (the space of all essentially bounded functions on $\mathbb{T}$), we follow the standard notation of the Toeplitz operator as $T_{\varphi}$, where
$T_{\varphi}f=P_{H^{2}}(\varphi f),$
for all $f\in H^{2}$, and $P_{H^{2}}$ denotes the orthogonal projection (or, the Szegö projection) from $L^{2}$ (the space of all square-integrable functions on $\mathbb{T}$) onto $H^{2}$. Pick an arbitrary orthonormal basis $\{f_{i}\}_{i=1}^{n}$ of $\mathcal{K}_{B_{n}}$, and consider the finite-rank perturbation of $T_{\varphi}$ as $T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i}$. In Theorem 7.1, we prove that
$\ker(T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i}),$
is nearly $T^{*}_{z,B_{n}}$-invariant subspace. We are of the opinion that this space is significant. We refer to [3, 7] for information regarding the kernels of Toeplitz operators and nearly $T_{z}^{*}$-invariant subspaces.
We also consider representations of invariant subspaces of finite-rank perturbations of $T_{z}$. We recall a general fact from elementary functional analysis. Let $\mathcal{S}$ be a closed subspace of a Hilbert space $\mathcal{H}$ and let $T\in\mathcal{B}(\mathcal{H})$. Then $\mathcal{S}$ is invariant under $T$ if and only if $\mathcal{S}^{\perp}$ is invariant under $T^{*}$. In particular, representations of $T$-invariant subspaces also yield representations of $T^{*}$-invariant subspaces. Nevertheless, the explicit feature of the invariant property (if any) may not be maintained when representations are transmitted from one to another, particularly in a concrete situation such as ours. Here, we are talking about the representation of invariant subspaces of finite-rank perturbations of the forward shift $T_{z}$, as we are already aware of the same for finite-rank perturbations of $T_{z}^{*}$.
We identify a class of perturbations of $T_{z}$ for which one can say more concretely about the lattice of invariant subspaces (see Theorem 4.1):
@@
Let $\mathcal{M}$ be an invariant subspace of $T_{z}-\sum_{i=1}^{m}v_{i}\otimes u_{i}$. Suppose
and let $p=\dim\mathcal{L}$. If $\{\varphi_{i}\}_{i=1}^{p}\subseteq H^{\infty}$ is an orthonormal basis for $\mathcal{L}$, then there exists an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ for some $p^{\prime}\leq p+1$ such that
@@
Some of the results stated above are merely one-directional. This is also in keeping with Hitt and Sarason’s perspectives on representations of nearly $T_{z}^{*}$-invariant subspaces. In some of the instances, we present a converse directional result that is restricted to a particular class of finite-rank perturbations. These perturbations are motivated by Sarason’s rank-one perturbation, as pointed out in (1.2). and typically they look like
$T_{z}^{*}-\sum_{i=1}^{n}T_{z}^{*}f_{i}\otimes f_{i}.$ (1.6)
@@
The remaining sections of the paper are structured in the following manner: Section 2 is the heart of the paper, which gives invariant subspaces of finite rank perturbations of Toeplitz operators with inner symbols vanishing at 00. In Section 3, we apply the results of the preceding section to the particular situation of the forward shift operator. We additionally take into account the Sarason-type perturbations. Section 4 deals with invariant subspaces of certain finite-rank perturbations of $T_{z}$. In Section 5, we continue our approach to finite-rank perturbations and determine the lattice of almost invariant subspaces of $T_{z}^{*}$. Section 6 introduces the concept of nearly $T^{*}_{z,B_{n}}$-invariant subspaces and presents an invariant subspace theorem. The final section, Section 7, is devoted to examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces. These subspaces are precisely the kernels of certain finite-rank perturbations of Toeplitz operators.

### v2

@@ Abstract.
We represent closed subspaces of the Hardy space that are invariant under finite-rank perturbations of the backward shift. We apply this to classify almost invariant subspaces of the backward shift and represent a more refined version of nearly invariant subspaces. Kernels of certain perturbed Toeplitz operators are examples of the newly introduced nearly invariant subspaces.
@@
Hardy space, invariant subspaces, shift operator, nearly invariant subspaces, model spaces, almost invariant subspaces
@@ 1. Introduction
The famous “invariant subspace problem” calls for finding a nontrivial closed subspace that is invariant under a bounded linear operator on a Hilbert space. This problem is still open. It is even worse that we do not know how to solve this problem for a general rank-one perturbation of a diagonal operator [6]. In the same vein, the lattice of invariant subspaces of a simple composition operator is unknown, and the issue is inextricably linked to the original invariant subspace problem [15]. Thus, the description of the lattices of invariant subspaces of common operators, or even the mere existence of invariant subspaces, is recognized as a complex problem. In particular, it is known to be a difficult challenge to find the lattice of invariant subspaces of finite-rank perturbations of well-known operators. In this paper, we will address this issue in relation to the well-known backward shift operator.
More specifically, in this paper, we bring together three problems concerning invariant subspaces of operators related to the backward shift operator $T_{z}^{*}$ acting on the Hardy space $H^{2}$. Recall the standard notation that $T_{z}$ is the multiplication operator by the coordinate function $z$; alternatively, the Toeplitz operator with the symbol $z$; and $H^{2}$ is the Hilbert space of all square summable analytic functions on the open unit disc $\mathbb{D}=\{z\in\mathbb{C}:|z|<1\}$ (see (2.3) for more details). Therefore
@@
for all $f\in H^{2}$. Strategically, we first represent closed invariant subspaces of finite-rank perturbations of $T_{z}^{*}$, and then we use that technique to classify almost invariant subspaces for $T_{z}^{*}$ and represent nearly invariant subspaces of $T_{z}^{*}$ (in a larger sense).
Before moving on, it is worth remarking that the results and approach of this paper are an enhancement of the theory that was presented by Hayashi [10] and Hitt [11], which was further fine-tuned by Sarason [18]. In fact, the idea we put forward for full-length parametrizations of invariant subspaces for all finite-rank perturbations of $T_{z}^{*}$ has been spurred by Sarason’s insight regarding the invariance property of a particular rank-one perturbation of $T_{z}^{*}$ [18, page 482] (also see (1.2) for more details). We have subsequently laid out the road map to solve the other two problems.
@@
for all $f\in H^{2}$. The term rank-one is used to describe operators whose ranges are one-dimensional. Similarly, rank-$m$ operators are those whose ranges are $m$-dimensional. They can be represented as:
@@
where $\{u_{i}\}_{i=1}^{m}$ and $\{v_{i}\}_{i=1}^{m}$ are orthonormal sets of functions in $H^{2}$. For the remainder of this section, we fix these orthonormal sets. We are interested in the invariant subspaces of a general finite-rank perturbation
@@
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under $X$, then there exist a natural number $p^{\prime}\leq p+1$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
@@
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F_{0}F+f_{0}\|^{2}=\|F\|^{2}+\|f_{0}\|^{2}\}.$
In the given context, $H^{2}_{\mathbb{C}^{p}}$ represents the $\mathbb{C}^{p}$-valued Hardy space, which may also be seen as $p$-copies of $H^{2}$. Given Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$, we denote $\mathcal{B}(\mathcal{H},\mathcal{K})$ (and $\mathcal{B}(\mathcal{H})$ if $\mathcal{K}=\mathcal{H}$) as the space of all bounded linear operators from $\mathcal{H}$ to $\mathcal{K}$. The symbol $H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ denotes the set of $\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})$-valued bounded analytic functions on $\mathbb{D}$, where a function $\Phi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ is inner if
$\Phi(z)^{*}\Phi(z)=I_{\mathbb{C}^{p^{\prime}}},$
for a.e. $z\in\mathbb{T}$. Moreover
$[F_{0},1]_{1\times(p+1)}=[f_{1},\ldots,f_{p},1],$
is a row vector, and $\mathcal{K}_{\Psi}$’s elements are columns.
In fact, we prove Theorem 1.1 in a much higher generality (see Theorem 2.1): we simply replace the backward shift $T_{z}^{*}$ by $T_{\varphi}^{*}$, the adjoint of a Toeplitz operator, where the symbol $\varphi\in H^{\infty}$ is an inner function vanishing at the origin.
Before proceeding, we need the definition of nearly $T_{z}^{*}$-invariant subspaces [18]. A closed subspace $\mathcal{M}$ of $H^{2}$ is nearly $T_{z}^{*}$-invariant if
@@
for all $f\in\mathcal{M}\cap zH^{2}$. Now, we will provide a few comments regarding Theorem 1.1. First, this result deals with linear perturbations, a theory in and of itself that is well-known to be challenging (see the classic [12]). Second, this parameterizes all of the invariant subspaces of finite-rank perturbations of $T_{z}^{*}$, which was known to be a challenging goal. As far as we know, this is the first result in this direction. Third, we provide this result in the tradition of Hayashi, Hitt, and Sarason’s classical argumentation. Indeed, Sarason showed [18, page 488] that if $\mathcal{M}$ is a nearly $T_{z}^{*}$-invariant subspace then there is a unique function
@@
such that $\mathcal{M}$ is invariant under the rank-one perturbation $T_{z}^{*}-T_{z}^{*}g\otimes g$, that is
@@
Sarason’s observation partly motivated us to attempt full-length representations of invariant subspaces of finite-rank perturbations, leading to the formation of Theorem 1.1. At the very start of this section, this point was briefly mentioned. Finally, we recall Hitt’s theory [11, 18] on almost $T_{z}^{*}$-invariant subspaces. Every nearly $T_{z}^{*}$-invariant subspace $\mathcal{M}\subseteq H^{2}$ admits the representation
@@
where $g$ is as in (1.1). Undoubtedly, the result in Theorem 1.1 is in line with the same fundamentals. As previously said, in the present case, we have partially enacted the core ideas that were put forward by Hayashi, Hitt, and Sarason a few decades ago.
Now we turn to almost invariant subspaces for $T_{z}^{*}$. In the setting of the backward shift operator, this was introduced by Chalendar, Gallardo-Gutiérrez, and Partington [4]. Let $T\in\mathcal{B}(\mathcal{H})$. A closed subspace $\mathcal{M}\subseteq\mathcal{H}$ is almost invariant for $T$ if there exists a finite-dimensional subspace $\mathcal{F}\subseteq\mathcal{H}$ such that
@@
In this case, the defect of the space $\mathcal{M}$ is the lowest possible dimension of such a space $\mathcal{F}$. A special version of Theorem 5.1 suggests a link between almost invariant subspaces for $T_{z}^{*}$ and invariant subspaces of finite-rank perturbations of $T_{z}^{*}$:
@@
Let $\mathcal{M}$ be a closed subspace of $H^{2}$. If $\mathcal{M}$ is invariant under
@@
then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect at most $m$. Conversely, if $\mathcal{M}$ is almost invariant under $T_{z}^{*}$ with defect $m$, then $\mathcal{M}$ is invariant under the rank-$t$ perturbation
@@
where $\{f_{i}:1\leq i\leq m\}$ is an orthonormal basis of the $m$ dimensional defect space and $t\leq m$.
In particular, this and Theorem 1.1 yield a complete characterization of almost invariant subspaces for $T_{z}^{*}$ (see Corollary 5.2):
@@
Let $\mathcal{M}\subseteq H^{2}$ be a closed subspace. If $\mathcal{M}$ is almost invariant under $T_{z}^{*}$, then there exist a natural number $p^{\prime}\leq p+1$ and an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ such that
@@
where $F_{0}=[f_{1},\ldots,f_{p}]$, $\{f_{i}\}_{i=1}^{p}$ is an orthonormal basis for $\text{span}\{P_{\mathcal{M}}T_{z}u_{i}:i=1,\ldots,m\}$, and $\{u_{i}\}_{i=1}^{m}$ is an orthonormal basis for the defect space. Moreover
$\mathcal{K}_{\Psi}=\{(F,f_{0})\in H^{2}_{\mathbb{C}^{p}}\oplus H^{2}:F_{0}F+f_{0}\in\mathcal{M}\textit{ and }\|F_{0}F+f_{0}\|^{2}=\|{F}\|^{2}+\|{f_{0}}\|^{2}\},$ (1.4)
@@
for all $(k_{1},\ldots,k_{p+1})\in\mathcal{K}_{\Psi}$, defines a unitary operator $\mathcal{J}:\mathcal{K}_{\Psi}\to\mathcal{M}$. Conversely, if $\mathcal{M}$ has the representation (1.3) for some orthonormal set of functions $\{f_{i}\}_{i=1}^{p}$ in $\mathcal{M}$, an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ with $\mathcal{K}_{\Psi}$ as in (1.4), and a unitary $\mathcal{J}$ as in (1.5), then $\mathcal{M}$ is almost invariant under $T_{z}^{*}$.
It should be noted that Chalendar, Gallardo-Gutiérrez, and Partington were the first to classify almost invariant subspaces under $T_{z}^{*}$ [4, Corollary 3.4]. However, the above result differs significantly from [4], which compares nearly $T_{z}^{*}$-invariant subspaces of finite defect indices with almost invariant subspaces. Here, we connect them with finite-rank perturbations of $T_{z}^{*}$. This, in particular, yields more concrete representations of the model space $\mathcal{K}_{\Psi}$. See the end of Section 5 for more comments.
Now we turn to a refined version of nearly $T_{z}^{*}$-invariant subspaces. The motivation for this arises partly from the matching theory of finite-rank perturbations of $T_{\varphi}^{*}$ as well as the potential of the possible theory of kernels of finite-rank perturbations of Toeplitz operators, which we will elaborate on shortly. We start by setting up the notations. Let $\{z_{k}\}_{k=1}^{n}$ be a sequence of points (possibly repeated) in $\mathbb{D}$, and assume that
@@
Evidently, nearly $T^{*}_{z,B_{1}}$-invariant subspaces are precisely nearly $T_{z}^{*}$-invariant subspaces. The following provides a complete description of nearly $T^{*}_{z,B_{n}}$-invariant subspaces of $H^{2}$ (for more details, see Theorem 6.2):
@@
Let $\mathcal{M}$ be a nontrivial closed subspace of $H^{2}$. Then $\mathcal{M}$ is nearly $T^{*}_{z,B_{n}}$-invariant if and only if there exist a natural number $r^{\prime}\leq r\leq n$ and an inner function $\Theta\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{r^{\prime}},\mathbb{C}^{r})}$ with $\Theta(0)=0$ such that
@@
where $\{g_{i}\}_{i=1}^{r}$ is an orthonormal basis for $\mathcal{M}\ominus(\mathcal{M}\cap B_{n}{H}^{2})$, and
@@
Denote by $L^{\infty}$ the space of all essentially bounded functions on $\mathbb{T}$. For each $\varphi\in L^{\infty}$, we follow the standard notation of the Toeplitz operator as $T_{\varphi}$. Then
$T_{\varphi}f=P_{H^{2}}(\varphi f),$ (1.6)
for all $f\in H^{2}$, where $P_{H^{2}}$ denotes the orthogonal projection (or, the Szegö projection) from $L^{2}$ (the space of all square-integrable functions on $\mathbb{T}$) onto $H^{2}$. Kernels of Toeplitz operators are crucial in the context of examples of nearly $T_{z}^{*}$-invariant subspaces. We refer to [3, 13, 17] for information regarding the kernels of Toeplitz operators and nearly $T_{z}^{*}$-invariant subspaces of $H^{2}$, and [9] in the context of $H^{p}$-subspaces.
In a comparable manner, we provide examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces. Curiously, certain dedicated finite-rank perturbations of Toeplitz operators result in nontrivial examples of $T^{*}_{z,B_{n}}$-invariant subspaces. Pick an arbitrary orthonormal basis $\{f_{i}\}_{i=1}^{n}$ of $\mathcal{K}_{B_{n}}$. In Theorem 7.1, we prove that
$\ker\Big{(}T_{\varphi}+\sum_{i=1}^{n}f_{i}\otimes T_{z}^{*}f_{i}\Big{)},$
is nearly $T^{*}_{z,B_{n}}$-invariant subspace. We are of the opinion that this space is significant.
We remark that Liang and Partington [13] also examined finite-rank perturbations of Toeplitz operators, providing examples of subspaces that are nearly $T_{z}^{*}$-invariant but have finite defects. They impose restrictions on the Toeplitz operators while leaving the finite-rank operators unrestricted. In this context, we also refer to [5]. However, our perspectives and objectives in this paper are different.
Next, we consider representations of invariant subspaces of finite-rank perturbations of $T_{z}$. We recall a general fact from elementary functional analysis. Let $\mathcal{S}$ be a closed subspace of $\mathcal{H}$ and let $T\in\mathcal{B}(\mathcal{H})$. Then $\mathcal{S}$ is invariant under $T$ if and only if $\mathcal{S}^{\perp}$ is invariant under $T^{*}$. In particular, representations of $T$-invariant subspaces also yield representations of $T^{*}$-invariant subspaces. Nevertheless, the explicit feature of the invariant property (if any) may not be maintained when representations are transmitted from one to another, particularly in a concrete situation such as ours. Here, we are talking about the representation of invariant subspaces of finite-rank perturbations $T_{z}$, as we are already aware of the same for finite-rank perturbations of $T_{z}^{*}$.
We identify a class of perturbations of $T_{z}$ for which one can say more definitively about the lattices of invariant subspaces (see Theorem 4.1):
@@
Let $\mathcal{M}\subseteq H^{2}$ be a closed subspace. Suppose $\mathcal{M}$ is an invariant under $T_{z}-\sum_{i=1}^{m}v_{i}\otimes u_{i}$. Define
$\mathcal{L}:=span\{P_{\mathcal{M}^{\perp}}v_{1},\ldots,P_{\mathcal{M}^{\perp}}v_{m}\}.$
Assume that $\{\varphi_{i}\}_{i=1}^{p}\subseteq H^{\infty}$ is an orthonormal basis for $\mathcal{L}$. Then there exists an inner function $\Psi\in H^{\infty}_{\mathcal{B}(\mathbb{C}^{p^{\prime}},\mathbb{C}^{p+1})}$ for some $p^{\prime}\leq p+1$ such that
@@
Some of the results stated above are merely one-directional. This is also in keeping with Hitt and Sarason’s perspectives on representations of nearly $T_{z}^{*}$-invariant subspaces. In some of the instances, we present a converse directional result that is restricted to a particular class of finite-rank perturbations. These perturbations are motivated by Sarason’s rank-one perturbation, as pointed out in (1.2). Typically they look like
$T_{z}^{*}-\sum_{i=1}^{m}T_{z}^{*}f_{i}\otimes f_{i},\text{ or }T_{z}^{*}-\sum_{i=1}^{m}f_{i}\otimes T_{z}^{*}f_{i}.$ (1.7)
@@
The remaining sections of the paper are structured in the following manner: Section 2 is the heart of the paper, which gives invariant subspaces of finite-rank perturbations of $T_{\varphi}^{*}$ with inner symbols $\varphi$ vanishing at zero. In Section 3, we apply the results of the preceding section to the particular situation of the backward shift operator. We additionally take into account the Sarason-type perturbations. Section 4 deals with invariant subspaces of certain finite-rank perturbations of $T_{z}$. In Section 5, we determine the lattice of almost invariant subspaces of $T_{z}^{*}$. Section 6 introduces the concept of nearly $T^{*}_{z,B_{n}}$-invariant subspaces and presents an invariant subspace theorem. The final section, Section 7, is devoted to examples of nearly $T^{*}_{z,B_{n}}$-invariant subspaces. These subspaces are precisely the kernels of certain finite-rank perturbations of Toeplitz operators.